Determine the following relation is an equivalence relation or not. If the relation is an equivalence relation, describe the partition given by it. xRy if x>yx,y∈the set of all real numbers
1
Expert's answer
2020-06-08T21:03:21-0400

The relation is given by

xRy if x>y , x,y "\in" "\mathbb {R}" , set of reals

i.e. R = { (x,y) : x > y , x,y "\isin \mathbb {R}" }

To show equivalency of the relation..

"\mathbf {REFLEXIVITY}" :

For a real x, x "\ngtr" x

So (x,x) "\notin" R

Therefore R is not reflexive

"\mathbf {SYMMETRY}" :

Let (x,y) "\isin" R for x,y "\in \mathbb {R}"

So x > y

Obviously y "\ngtr" x

Therefore (y,x) "\notin" R

So R is not symmetric.

"\mathbf {TRANSITIVITY}" :

Let (x,y) "\in" R and (y,z) "\isin" R for x,y,z "\in \mathbb {R}"

So x > y and y > z

Obviously x > z

So R is transitive.


A relation be equivalence relation if it is reflexive, symmetric and transitive.

In this case R is transitive but neither reflexive nor symmetric.

So given R is not an equivalence relation.

As it is not equivalence relation , second part of the question is not applicable.